Given a body, of negligible weight, initially a point (20,30) in the SI convention on the x-y plane, is fired at 400m/s at an angle of 60degrees(to the vertical). What coordinates will the body assume after 30sec of travel

1
Expert's answer
2021-11-01T12:53:04-0400

The horizontal and vertical components of velocity:


"v_x=v\sin\theta,\\
v_y=v\cos\theta."

The x-coordinate after 30 s:


"x=x_0+v_xt=x_0+vt\sin\theta=10412\text{ m}."


The y-coordinate:


"y=y_0+v_yt-\frac{gt^2}2=y_0+vt\cos\theta-\frac{gt^2}2=1620\text{ m}."